OpenStudy (loser66):

a,b,c,d \(\in R\) ; x, y \(\in C^2\) find the condition of a, b, c, d to get the > 0

OpenStudy (loser66):

a,b,c,d are real \(x = (x_1,x_2)\); \(y =(y_1,y_2)\) ; the inner product defined \[<x,y>= ax_1\bar{y_1}+bx_2\bar{y_1}+cx_1\bar{y_2}+dx_2\bar{y_2}\] how to prove \(<x,x>\geq 0\) = iff x =0 Please, help

OpenStudy (loser66):

so far, I have \[<x,x>= ax_1\bar{x_1}+bx_2\bar{x_1}+cx_1\bar{x_2}+dx_2\bar{x_2}\] \[=a|x_1|^2 +bx_2\bar{x_1}+cx_1\bar{x_2} +d|x_2|^2\] and stuck

OpenStudy (phi):

which are you doing ? find the condition of a, b, c, d to get the <x,x>> 0 or how to prove <x,x>≥0 = iff x =0

OpenStudy (loser66):

the first one

OpenStudy (loser66):

the condition of a, b, c, d

OpenStudy (phi):

do we need <x,x> to be real ?

OpenStudy (loser66):

we need it complex to have conjugate of the term \(\bar x_1\) for example

OpenStudy (phi):

I mean what does it mean for a complex number C to be ≥ 0 ? as opposed to | C | ≥ 0

OpenStudy (loser66):

I don't get your question

OpenStudy (phi):

\[ =a|x_1|^2 +bx_2\bar{x_1}+cx_1\bar{x_2} +d|x_2|^2 \] this will be >0 if a and d >0 and b=c=0 that might be to constraining, but it works.

OpenStudy (loser66):

Is there any other option else?

OpenStudy (anonymous):

Wouldn't \(\bar{y_1}=y_1\)? If \(y\in\mathbb{C}\), then \(y_1,y_2\in\mathbb{R}\). Just wondering if you really mean to use the conjugate there.

OpenStudy (anonymous):

Also, @loser66 you mentioned that you had to prove \(\langle x,x\rangle \ge0\) iff \(x=0\). Do you mean \(\langle x,x\rangle=0\) iff \(x=0\)?

OpenStudy (phi):

we can write this problem as \[ u^T M u ≥ 0\] where M= [a b ; c d] that means we want M to be positive definite

OpenStudy (loser66):

No, no @SithsAndGiggles, x, y in C^2, it means x1, x2, y1, y2 are complex

OpenStudy (loser66):

This is the whole thing and what I post here is the last part where I stuck

OpenStudy (phi):

Are you sure b and c have to be real? I am thinking M has to be hermetian, i.e. M = conj( transpose (M))

OpenStudy (loser66):

when proving the defined equation is an inner product, in part 1, I have a, b,c, d must be real. That's why in part 3, I use that condition to figure out the other condition. brb

OpenStudy (phi):

I think you can show that a and d are real, and b = conj( c ) define <x,y> = y^H M x where H means complex conjugate transpose, and M = M^H we can show that <y*, x*> = <x,y>

OpenStudy (loser66):

let me try, thanks for the help :)

OpenStudy (phi):

something like this

OpenStudy (loser66):

yes, that 's what the first part I did and get the first condition is a, b, c, d are real

OpenStudy (phi):

yes for a and d real. b = conjugate of c (not necessarily real)